OpenStudy (anonymous):

solve the following linear equations by substituting: -4x+2y=-12 and 8x+4y=24

OpenStudy (anonymous):

4y=-8x+24 y=-2x+6 Now plug that into y in following equation -4x+2y=-12

OpenStudy (anonymous):

you can find x find setting y=2x-6 =-2x+6

OpenStudy (anonymous):

If you're finding the solution there's much more to do

OpenStudy (anonymous):

The solution to 2 crossing lines will be a point, (x,y)

OpenStudy (anonymous):

trying to find out if its consistent or inconsistent or dependant

OpenStudy (anonymous):

2x-6=-2x+6 4x-12 4x=12 x=3

OpenStudy (anonymous):

Now plug that into one of the equation to find y

OpenStudy (anonymous):

Inconsistent means no solution. Consistent means 1 solution. Dependant means infinite solutions (same line).